OpenStudy (anonymous):

A leaky valve on the water meter overcharges the residents for one gallon of water in every 2 1/2 months. The overcharged amount w varies directly with time t. a. Find the equation that models this direct variation. b. How many months it will take for the residents to be overcharged for 8 gallons of water?

OpenStudy (hba):

algb 2 righht

OpenStudy (anonymous):

yes!

OpenStudy (hba):

w=3/4t and 6 2/3 months option a

OpenStudy (anonymous):

thats answer A?

OpenStudy (hba):

yeah

OpenStudy (anonymous):

welll all my answer choices have 2/5 in them..

OpenStudy (anonymous):

w=2/5t ; 20 months

OpenStudy (anonymous):

w=5/2t; 20 months

OpenStudy (anonymous):

w=2/5t; 3 1/5 months